# Cost-aware Portfolios in a Large Universe of Assets

### v1

@@ 4.2 Russel 2000 Index
In this subsection we evaluate the performance of CAPE in the relatively small-cap stocks pool. We collected daily Russell 2000 Index composition stocks data for the holding period from 02-Jan-2017 to 23-Dec-2020. Our portfolio is rebalanced every 251 trading days, resulting in three decision days and four stages. The out-of-sample period extends from 01-Jan-2018 to 23-Dec-2020. We employ the same criterion of complete observations described in Section 4.1, resulting in a pool of 991 stocks. This scenario imposes great challenges for portfolio management in practice, however, it showcases the advantage and applicability of our proposed method. The performance of the various methods is reported in Tables 6 and 7, which correspond to proportional and quadratic transaction costs, respectively.
The performances of MV and CMV methods are notably poor, as demonstrated in Tables 6 and 7. The PMV method exhibits a slightly better overall Sharpe ratio during the S2 to S4 phases compared to the MV and CMV methods. In contrast, our CAPE method significantly outperforms the other approaches, achieving the highest overall Sharpe ratio. The turnover ratio of CAPE is notably the smallest among all actively managed portfolios. Consistent with our findings in the S&P 500 large-cap stocks, the CAPE method demonstrates robustness to variations in transaction cost functions.
Table 6: Russell 2000 quadratic transaction costs, the unit rebalancing cost of the portfolio $\beta=0.15$ and $\gamma=1/3$.
@@
Table 7: Russell 2000 proportional transaction costs, the unit rebalancing cost of the portfolio $\alpha=0.001$ and $\gamma=1/3$.
@@ 5 Conclusion
In this paper, we propose a novel approach to integrate the transaction cost within the optimization problem when the number of assets is larger than the sample size. By using LLA algorithm, we prove the sign consistency and provide the $\ell_{\infty}$ bound of our proposed CAPE method for both proportional and quadratic transaction costs.
In addition to the sound statistical properties, CAPE demonstrates exceptional performance in practical applications. Through extensive simulations and applications to real data we demonstrate the superior performance and broad applicability of our method across diverse tasks.

### v2

@@ 4.2 Russel 2000 Index
Next, we report the results for the Russell 2000 dataset, following the same setup as in Section 4.1. The corresponding performance metrics are presented in Tables 6 and 7.
Table 6: Russell 2000 with quadratic transaction cost.
@@
Table 7: Russell 2000 with proportional transaction cost.
@@
Despite the higher volatility and liquidity challenges inherent in small-cap stocks, the CAPE-S method maintains robust performance under both quadratic and proportional transaction cost settings. Tables 6 and 7 report detailed results under the two cost regimes.
  • Stage S1 (2018) was a period of elevated volatility and broad market drawdown. The Russell 2000 Index, which is generally more sensitive to macroeconomic uncertainty, suffered larger losses than the S&P 500. In this challenging environment, CAPE-S was the only strategy to achieve positive returns, with modest gains accompanied by low turnover (approximately 1.3%) and low leverage (around 0.15), highlighting its defensiveness and cost-efficiency.
  • Stage S2 (2019) was characterized by a broad market rebound, driven by easing monetary policy and reduced trade tensions. During this recovery phase, CAPE-S achieved the highest Sharpe ratio (2.745) among all strategies, while maintaining the lowest transaction cost (0.280%) and turnover (approximately 1.52%), far below the MV strategy’s cost of 23.272% and turnover of 34.829%. This demonstrates that CAPE-S effectively balances risk and return without relying on aggressive rebalancing.
  • Stage S3 (2020) encompassed one of the most turbulent and rapidly rebounding market periods in recent history due to the COVID-19 pandemic. The small-cap segment staged a dramatic recovery following the March 2020 crash, benefiting disproportionately from fiscal stimulus and risk-on sentiment. CAPE-S captured substantial gains (approximately 177.99%) while preserving a favorable risk profile, with low leverage (0.973) and turnover (1.60%), outperforming all alternatives in terms of Sharpe ratio (1.710).
Overall, the MV and CMV strategies performed poorly, particularly under proportional cost settings, due to high turnover and leverage. The PMV method showed some improvement but still lagged behind CAPE-S. Across all three stages, CAPE-S consistently achieved the highest overall Sharpe ratio (1.170 for quadratic costs and 1.158 for proportional costs) while maintaining the lowest turnover among actively managed portfolios. These findings confirm CAPE-S’s adaptability and robustness, particularly in the more volatile and cost-sensitive small-cap universe, and underscore its practical relevance for real-world portfolio management.
@@ 5 Conclusion
In this paper, we propose a novel approach to integrate the transaction costs within the optimization problem when the number of assets is larger than the sample size. By using LLA algorithm, we prove the sign consistency and provide the $\ell_{\infty}$ bound of our proposed CAPE-S method for both proportional and quadratic transaction costs.
In addition to the sound statistical properties, CAPE-S demonstrates exceptional performance in practical applications. Through extensive simulations and applications to real data, we demonstrate the superior performance and broad applicability of our method across diverse tasks.